# Representational Stability of Truth in Large Language Models

### v2

@@ Appendix E Exploring Additional Probes
We repeated the label perturbation experiments using the Mean Difference probe proposed by Marks and Tegmark [marks2310geometry]. It estimates a “truth direction” by taking the vector difference between the mean activation of True statements and that of False statements, optionally scaled by the inverse covariance matrix of the data. This approach is inherently sensitive to differences in the centroids and covariance structure of the data, which leads to strong instability in the learned decision boundary when Neither statements are included alongside true and false examples. The Mean Difference probe show considerably greater variability across LLMs than sAwMIL (Fig. A17). While sAwMIL yields consistent decision boundary rotation corresponding to specific perturbations, particularly the Synthetic perturbation, the Mean Difference probe exhibits near-orthogonal boundary shifts for certain LLMs regardless of perturbation. In addition, Table A4 shows that, unlike with sAwMIL, the Fictional perturbation produces the largest number of prediction flips across all three datasets, and the Word Definitions dataset exhibits the fewest total flips. We interpret these discrepancies as artifacts of the Mean Difference probe’s reliance on dataset centroids: when statement activations are well separated, as with Fictional statements, class-label perturbations can induce disproportionately large changes in the estimated decision boundary. This instability reflects probe sensitivity rather than genuine representational instability in the LLMs. Accordingly, the Mean Difference probe is less well suited for quantifying representational stability than sAwMIL.
@@
Table A4: Flipped predictions under label perturbations for the Mean Difference Probe. Counts (and percentages) of predictions that remain stable or flip between True and Not True across Synthetic, Fictional, Fictional(T), and Noise perturbations for each dataset. Word Definitions is the most stable, followed by City Locations and Medical Definitions. The Fictional perturbation leads to the most instability.
Dataset Perturbation True to True Not True to Not True Not True to True True to Not True Total Flips
City Locations Synthetic $9724$ $(97.2)$ $167$ $(1.7)$ $63$ $(0.6)$ $46$ $(0.5)$ $109$ $(1.1)$
Fictional $7386$ $(73.9)$ $221$ $(2.2)$ 99 $(0.1)$ $2384$ $(23.8)$ $\mathbf{2393}$ $\mathbf{(23.9)}$
Fictional (T) $9680$ $(96.8)$ $184$ $(1.8)$ $46$ $(0.5)$ $90$ $(0.9)$ $136$ $(1.4)$
Noise $9653$ $(96.5)$ $201$ $(2.0)$ $29$ $(0.3)$ $117$ $(1.2)$ $146$ $(1.5)$
Medical Locations Synthetic $9457$ $(86.8)$ $894$ $(8.2)$ $221$ $(2.0)$ $324$ $(3.0)$ $545$ $(5.0)$
Fictional $4694$ $(43.1)$ $977$ $(9.0)$ $138$ $(1.3)$ $5087$ $(46.7)$ $\mathbf{5225}$ $\mathbf{(48.0)}$
Fictional (T) $9718$ $(89.2)$ $1069$ $(9.8)$ $46$ $(0.4)$ $63$ $(0.6)$ $109$ $(1.0)$
Noise $8955$ $(82.1)$ $1026$ $(9.4)$ $89$ $(0.8)$ $826$ $(7.6)$ $915$ $(8.4)$
Word Definitions Synthetic $8468$ $(85.6)$ $500$ $(5.1)$ $695$ $(7.0)$ $225$ $(2.3)$ $920$ $(9.3)$
Fictional $7035$ $(71.1)$ $1175$ $(11.9)$ 2020 $(0.2)$ $1658$ $(16.8)$ $\mathbf{1678}$ $\mathbf{(17.0)}$
Fictional (T) $8282$ $(83.8)$ $1083$ $(11.0)$ $112$ $(1.1)$ $411$ $(4.2)$ $523$ $(5.3)$
Noise $8208$ $(83.0)$ $1178$ $(11.9)$ $17$ $(0.2)$ $485$ $(4.9)$ $502$ $(5.1)$

### v3

@@ Appendix E Exploring the Mean Difference Probe
We repeated the representation-based perturbation experiments using the Mean Difference probe proposed by Marks and Tegmark [20] to supplement the sAwMIL results. Mean Difference estimates a “truth direction” by taking the vector difference between the mean activation of True statements and that of False statements, optionally scaled by the inverse covariance matrix of the data. This approach is inherently sensitive to differences in the centroids and covariance structure of the data, which leads to strong instability in the learned decision boundary when Neither statements are included alongside True and False examples. The Mean Difference probe shows considerably greater variability across LLMs than sAwMIL (Fig. A25). While sAwMIL yields consistent decision boundary rotation corresponding to specific perturbations (see Fig. A17, particularly the Synthetic perturbation), the Mean Difference probe exhibits near-orthogonal boundary shifts for certain LLMs regardless of perturbation. In addition, Table A24 shows that, unlike with sAwMIL, the Fictional perturbation produces the largest epistemic retractions across all three datasets, and the Word Definitions dataset exhibits the fewest total retractions. We note, however, that the Synthetic perturbation still produces the most epistemic expansions across domains, consistent with the sAwMIL results. We interpret these discrepancies as artifacts of the Mean Difference probe’s reliance on dataset centroids: when statement activations are well separated, as with Fictional statements, class-label perturbations can induce disproportionately large changes in the estimated decision boundary. This instability reflects probe sensitivity rather than genuine representational instability in the LLMs. Accordingly, the Mean Difference probe is less well suited for quantifying stability within P-StaT than sAwMIL.
@@
Figure A24: Epistemic expansions $\mathcal{E}$ and retractions $\mathcal{R}$ under probing label perturbations for the Mean Difference Probe. Counts (and percentages) of beliefs that remain stable or lead to expansions $\mathcal{E}$ and retractions $\mathcal{R}$ across Synthetic (yellow), Fictional (gray), Fictional(T) (blue), and Noise (red) perturbations for each dataset. While the Fictional perturbation induces the most epistemic retractions $\mathcal{R}$, the Synthetic perturbation induces the most epistemic expansions $\mathcal{E}$.